Grade 12th passAIIMS MBBS

Two simple pendulums of length 1 m and 16 m are in phase at the mean position at a certain instant of time. If T is the time period of shorter pendulum then the minimum time after which they will again in phase is

Aun
8 Years agoGrade 12th pass

1 Answer

roopali upadhha
8 Years ago
Since the time period of shorter pendulum is T.Therefore time period of pendulum of length 16is 4T. let t be time when they are in phase.let n be no of oscillation of slow(length 16) and (n+1) that of fast(length=1) t=n×4T=(n+1)T4n=n+1n=1/3 for 16 cm lengthfor shorter pendulum n+1=4/3 so4/3 ×T is minimum time after which they will be in phase again